OpenStudy (anonymous):

Rafael has 1 black, 3 red, 7 green, 2 blue, and 2 yellow candies in his pocket. What is the probability he will pull a red one out of his pocket? He has 15 candies in all.

OpenStudy (anonymous):

1/5 because he has 3 red and a total of 15. 3/15=1/5

OpenStudy (anonymous):

(total number of wanted color)/ (total number of all)

OpenStudy (apoorvk):

1/5... (probab=(no. of favourable outcomes)/(no.of favourable outcomes)

OpenStudy (apoorvk):

Ooops... probab=(no. of favourable outcomes)/(no.of "POSSIBLE" outcomes)

OpenStudy (anonymous):

1/15 because the red candies are 1 of the choices out of 15 candies idk for positive but i think its correct

OpenStudy (anonymous):

*for positive